Find the ratio of: 
9 m to 27 cm
Find the ratio of: 
9 m to 27 cm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that,
1 m = 100 cm
Then,
9 m = 9 × 100 = 900 cm
Now we have to find the ratio,
= 900/27 ... [∵divide both by 9]
= 100/3
So, the required ratio is 100: 3

Thanks